# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v3

@@ 5 Experiments
In this section, we present two main sets of experiments: (i) experiments with white-box access to the conditional probability distributions for token generation to examine our theoretical results, and (ii) experiments that relax the white-box access assumption. In the black-box setting, we also examine robustness to decoding strategies and adversarial attacks.
Figure 2: Overview of the experiments
Lastly, we provide experiments to demonstrate that the log-perplexity under the evaluator model and the average cross-entropy between the generative and the evaluator model converge.
5.1 White-box detection
Datasets. We use the datasets used in Mitchell et al. (2023) for our experiments. That includes the news articles from the XSum dataset (Narayan et al. 2018), prompted stories from Reddit’s WritingPrompts dataset (Fan et al. 2018), and Wikipedia paragraphs from the SQuAD contexts (Rajpurkar 2016) that contain human-written questions and machine-generated answers from passages from Wikipedia articles. Each experiment uses 150 to 500 examples for evaluation. We prompt the first 30 tokens of the actual text to the LLM to generate machine-generated text and obtain the generated text. For QA datasets, we prompt the question. In all experiments, a positive signal is that the focal LLM generates text, and all experiments use an equal number of positive and negative examples. The language models we examine for whitebox detection are LLAMA 3 (8B parameters), GPT-NEOX Erebus (20B parameters), and QWEN (32B parameters). We use AUROC and TPR @ FPR=1%, which follows the recent research that uses this evaluation paradigm (Krishna et al. 2024; Hans et al. 2024; Soto et al. 2024) as our performance metrics.
AUROC. The Area Under the Receiver Operating Characteristic (ROC) Curve is a standard widely adopted detection performance measure, e.g., Verma et al. 2023; Mitchell et al. 2023, and we use this metric consistent with the literature. The ROC curve plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at all possible classification thresholds, showing the tradeoff between correctly detecting LLM-generated text and mistakenly flagging human-written text.
TPR @ FPR=1%. Results of experiments in the literature, e.g., in Dugan et al. 2024, suggest that the false positive rates of detectors can be dangerously high, which makes them unfit for regulatory actions. To show the validity of our theoretical results, we examine the accuracy captured by the TPR while maintaining the low FPR.
Benchmarks. We compare the performance of our method with various existing zero-shot detection methods that similarly leverage the predicted token-wise conditional distributions of the source model. The benchmarks we use apply detection tests based on token log-probabilities, token (log-)ranks, likelihood curvature, or (a measure of) predictive entropy (Gehrmann et al. 2019; Solaiman et al. 2019; Ippolito et al. 2019; Mitchell et al. 2023; Hans et al. 2024). Detection based on token log-probabilities uses the source model’s average token-wise log-probability. Detection based on token (log-)rank uses the average observed (log-)rank of the tokens in the passage based on the model’s conditional distributions. The entropy-based detection benchmark inspired by Gehrmann et al. 2019 and applied by Mitchell et al. 2023 hypothesizes that machine-generated texts are more "in-distribution" for the model, leading to lower entropy predictive distributions. Detection based on Mitchell et al. 2023’s framework states that a machine-generated text is sampled mostly at the mode of its text distribution, while human texts might lay anywhere on the text distribution. Lastly, Hans et al. (2024) applies a normalized measure of perplexity by dividing log-perplexity by average cross-entropy.
5.2 White-box detection results
AUROC. In these experiments, model samples are generated by sampling from the raw conditional distribution with temperature 1.
XSum SQuAD WritingPrompts
$\log p(x)$ 0.99* 0.84 0.99* 0.91 0.75 0.60 1.00 0.95 1.00*
Rank 0.75 0.69 0.69 0.71 0.71 0.55 0.82 0.80 0.81
LogRank 0.99* 0.87 0.99* 0.93* 0.81 0.62 1.00 0.97* 1.00*
Entropy 0.39 0.70 0.40 0.37 0.66 0.70* 0.04 0.36 0.02
DetectGPT 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Binoculars 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98* 1.00 1.00
Diff (zero-shot) 0.00 0.04 0.01 0.06 0.18 0.27 -0.02 0.03 0.00
Roberta (base) 0.98* 0.95 0.92* 0.97* 0.92 0.69* 0.97* 0.95* 0.74*
Roberta (large) 0.98* 0.98* 0.92* 0.95 0.93* 0.68 0.96 0.93 0.65
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98 1.00 1.00
Diff (supervised) 0.01 0.01 0.08 0.02 0.06 0.28 0.01 0.05 0.26
Table 1: AUROC for detecting samples from the given models on the datasets using our and other baseline methods. The best AUROC values are in bold, and the second-best values are marked with an asterisk (*). The rows Diff(zero-shot) and Diff(supervised) show our AUROC improvement over the strongest zero-shot and supervised baseline methods.
Our method achieves its highest improvement in detecting LLM-generated text for SQuAD, increasing AUROC from 0.04 to 0.27. It also improves detection for XSum (0.015 AUROC) and WritingPrompts (0.025 AUROC), though with a smaller margin. An explanation for the weaker performance of existing detectors on SQuAD is its higher entropy variance (0.128) compared to XSum (0.045) and WritingPrompts (0.074). Since our method relies on the difference between log-perplexity and empirical entropy, it performs more effectively when entropy variance is high while aligning with likelihood-based methods when entropy is less variable. Finally, our method significantly outperforms existing likelihood-based approaches in detecting Qwen-generated text for SQuAD.
TPR @ FPR=1%. In Table 2, we show the accuracy captured by TPR, while the FPR is maintained low (1%).
@@ 5.2 White-box detection results
The experiment’s results show the notable superiority of our method in fact-based datasets (e.g., SQuAD), where the human-written and LLM-generated texts tend to be similar, resulting in lower performance of conventional detection methods. Specifically, we observe the maximum gap between ours and the second-best methods in the SQuAD dataset with QWEN 32B as the evaluator. Even when the FPR is held low, the human and LLM texts are more distinguishable in datasets with less fact-based and more creative content. Also, smaller language models (e.g., LLAMA 2 8B) are easier to detect.
5.3 Black-box detection
Datasets. We use the datasets in Dugan et al. 2024 that consist of data in 8 domains of abstracts, books, news, poetry, recipes, Reddit, reviews, and Wikipedia. The non-adversarial dataset consists of 14,971 human-written from publicly-available pre-2022 datasets and 509,014 LLM-generated documents for a total of 6,287,820 texts when including adversarial attacks.
@@ 5.3 Black-box detection
Benchmarks. We compare the performance of our method with the detectors in Dugan et al. 2024 that include neural detectors, zero-shot detectors, and commercial detectors. The neural detectors are RoBERTa-B (GPT-2), RoBERTa-L (GPT-2), RoBERTa-B (ChatGPT), and RADAR. The zero-shot detectors are GLTR, Fast DetectGPT, Binoculars, and LLMDet. The commercial detectors are GPTZero, Originality, Winston, and ZeroGPT. In this section, we do not have white-box access to any of the detectors, and we use tiiuae/falcon-rw-1B as our scoring model.
Repetition penalty. This penalty down-weighs the probability of tokens that previously appeared in the context window by some multiplicative factor $\theta$, resulting in less repetitive output. Following Keskar et al. 2019, we use $\theta=1.2$ for our experiments.
Adversarial attacks. When selecting adversarial attacks, consistent with Dugan et al. 2024, we assume that the adversary uses predefined transformations or heuristics that degrade typical detection cues, different from the gradient-based methods. Accordingly, we include the following 8 black-box, query-free attacks:
  1. 1.
    Alternative Spelling (AS): uses British spelling instead of the American spelling.
  2. 2.
    Article Deletion (AD): deletes "the", "a", "an."
  3. 3.
[…]
  4. 4.
    Number Swap (NS): randomly shuffles number digits.
  5. 5.
    Misspelling (MS): inserts common misspellings.
  6. 6.
    Synonym Swap (SS): swaps tokens with highly similar BERT (Devlin et al. (2019)) candidate tokens.
  7. 7.
    Upper Lower Swap (ULS): swaps the case of words.
  8. 8.
    Whitespace Addition (WSA): adds spaces between characters.
Note that two other adversarial attacks, i.e., Homoglyph and Zero-width space, are possible. Homoglyph swaps characters in a text with visually similar yet technically different characters, often from other alphabets or Unicode blocks. Zero-width space inserts a Unicode character (often U+200B) that does not occupy visible space. Dugan et al. (2024)’s dataset for these attacks was incomplete, not allowing us to provide results for these attacks. However, these two attacks are easy to remove with the following pre-processing:
  • Removing or normalizing invisible characters like zero-width spaces
  • Mapping suspicious non-ASCII characters to standard ASCII or a canonical form
5.4 Black-box detection results
Accuracy at FPR=5%. Consistent with Dugan et al. (2024), we evaluate accuracy by TPR while fixing FPR at 5%. In the experiments, the open-source chat models include Llama-c, Mistral-c, and MPT-c. The open-source non-chat models include Mistral, MPT, and GPT2. The closed-sourced chat models include c-GPT, GPT4, and Cohere, and the closed-sourced non-chat models include Cohere and GPT3. For the repetition penalty, we set the multiplicative factor as $\theta=1.2$.
@@ 5.4 Black-box detection results
Rep. Penalty? × × × ×
GPTZero 98.4 82.5 9.4 4.8 88.5 53.4
Originality 97.7 72.5 89.0 51.2 89.0 85.4
Winston 96.6 78.3 29.5 11.3 93.7 68.1
ZeroGPT(*) 90.5 54.9 16.0 0.3 65.8 72.7
R-B GPT2 77.9 26.2 60.5 35.4 41.7 52.5
R-L GPT2 71.4 19.5 67.2 53.4 34.7 48.6
R-B CGPT 75.0 39.3 14.9 1.7 38.1 39.0
RADAR 85.6 66.4 48.3 31.8 75.3 67.7
GLTR 83.9 38.3 44.5 0.5 54.3 63.7
F-DetectGPT 96.2 40.5 79.7 0.6 74.1 86.3
LLMDet 47.5 16.5 38.4 3.7 18.5 32.9
Binoculars 99.7 60.6 72.4 0.6 92.1 95.0
Ours 97.6 60.7 51.7 71.2 68.8 86.5
Table 3: Accuracy Score at FPR=5% for all detectors across model groups and sampling strategies. Asterisks (*) indicate that the detector was unable to achieve the target FPR. For accuracy of 67%-100%, we highlight green, 33%-66% we highlight yellow, and for 0%-32% we highlight red.
For open-source chat models without a repetition penalty, our method achieves an accuracy of 97.6%, indicating that when text maintains its natural stylistic and structural cues, our detector can accurately flag LLM-generated content. For open-source non-chat models, the addition of a repetition penalty improves performance, with accuracy rising from 51.7% without the penalty to 71.2% with the penalty. In the closed-source setting, our detector again performs robustly—68.8% for chat models and a high 86.5% for non-chat models, demonstrating our method’s capacity to generalize across different text styles. Overall, these findings indicate that our detector is particularly strong when the underlying text generation preserves the natural language features that our method is designed to leverage, and that applying a repetition penalty can further enhance accuracy in contexts where repetitive patterns might otherwise obscure these signals.
Adversarial attacks
None AS AD IP NS MS SYN ULS WSA
RoB-B GPT2 59.1 55.6 37.1 56.9 55.9 43.8 71.5 18.8 45.2
RoB-L GPT2 56.7 52.4 33.2 55.1 51.7 39.5 79.4 19.3 40.1
RoB-B CGPT 44.8 43.3 38.0 5.2 44.3 42.1 39.6 31.7 0.1
RADAR 70.9 70.8 67.9 73.7 71.0 69.5 67.5 70.4 66.1
GPTZero 66.5 64.9 61.0 66.2 65.8 65.1 61.0 56.5 66.2
ZeroGPT 65.5 65.4 59.7 64.9 64.7 64.7 18.8 54.5 64.2
Originality 85.0 83.6 71.4 85.1 86.0 78.6 96.5 75.8 84.9
Winston 71.0 68.9 66.9 69.8 69.0 67.5 63.6 56.8 46.8
GLTR 62.6 61.2 52.1 61.4 59.9 59.8 31.2 48.1 45.8
F-DGPT 73.6 71.6 64.7 72.0 68.2 70.7 34.0 60.4 64.4
LLMDet 35.0 33.9 27.4 27.2 33.8 32.7 27.3 23.4 4.4
Binoculars 79.6 78.2 74.3 71.7 77.1 78.0 43.5 73.8 70.1
Ours 84.3 79.9 73.1 80.7 73.3 78.1 48.0 72.9 57.1
@@
Table 4: Accuracy Score at FPR=5% for all detectors across different adversarial attacks. Abbrevations are: AS: Alternative Spelling, AD: Article Deletion, IP: Insert Paragraphs, NS: Number Swap, MS: Misspelling, SYN: Synonym Swap, ULS: Upper Lower Swap, WSA: Whitespace Addition. Cell colors: Red < 33, Yellow 33–66, Green $\geq$ 67.
Our detector maintains relatively high accuracy across multiple adversarial scenarios. In the baseline “None” condition, it records an accuracy of 84.3%. Under attacks such as Alternative Spelling (79.9%), Article Deletion (73.1%), Insert Paragraphs (80.7%), Number Swap (73.3%), Misspelling (78.1%), and Upper–Lower Swap (72.9%), the accuracy remains in the 70–80% range. However, similar to other zero-shot detectors, our performance is reduced under Synonym Swap (48.0%) and Whitespace Addition (57.1%). This detection performance drop is predictable because zero-shot detection methods often rely on fixed token distributions. Synonym substitutions replace surface-level tokens with semantically equivalent but lexically distinct items, thereby shifting the token distribution away from the detector models’ reference distribution. Whitespace manipulations alter standard sub-word segmentation by inserting or removing token boundaries, creating a discrepancy between the text’s actual token frequency and what the detector expects.
5.5 Convergence results
We conduct our numerical analysis for convergence on the following pre-trained language models: GPT-2 small, GPT-2 medium, GPT-2 large, GPT-2 XL, and GPT-Neo. Through experiments with different generative and evaluator models, we examine whether the log-perplexity of a short portion of text converges to the average cross-entropy. Our experiments measure these values across generated text and analyze their performance over different configurations. Our setup includes generating tokens with pre-trained models and recording each token’s selection probability and calculated metrics.
Same generative and evaluator model. In the first set of experiments, we employ GPT-2 to generate a series of $100$ tokens, beginning with the fixed prompt "Jack". We use the model’s conditional probability distribution for each token generation step to sample the next token. Note that for the white-box model of GPT-2, probability distributions are accessible. We calculate each generated token’s empirical entropy and log-perplexity and repeat this process for comparisons. We use Softmax-normalized probabilities to select the next token and store the generated token and its probability distribution. For each sub-string of length $N$ starting from the first token in the generated sequence, we compute the log-perplexity $l_{A}(\mathbf{Y}_{N})$, and the empirical entropy $h_{N}(A,A)(\mathbf{Y}_{N})$. The results are shown in Figures (3(a)-3(d)). We consistently observe that the numerical results confirm Theorem 1 that the log-perplexity converges to the average entropy when the generative and evaluator models are the same.
(a) GPT-2 small
(b) GPT-2 medium
(c) GPT-2 large
(d) GPT-2 XL
Figure 3: Generated and evaluated by the same model
Different generative and evaluator models. To extend our numerical analysis to the case with different generative and evaluator models, we generate a string using the following generative models: GPT-2 medium, GPT-2 large, and GPT-2 XL. Then, we calculate the log-perplexity of these strings using the evaluator model GPT2-small. We calculate the cross-entropy of the strings under each generative model and the evaluator model (GPT2- small). The results are shown in Figures (4(a)-4(c)). Results in these figures confirm Theorem 2. In particular, we observe that when the evaluator and generative models are different, the log-perplexity of the string converges to the average cross-entropy of the string under generative and evaluator models.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XL
Figure 4: Generated by GPT-2 small and evaluated by different models

### v4

@@ 5 Experiments
This section presents three sets of experiments: Subsection 5.1 examines Proposition 1, and Subsection 5.2 examines Proposition 2 where we use human text to represent the null hypothesis (not LLM $A$). Subsection 5.3, using a dataset with approximately 6M texts, empirically examines the robustness of Proposition 2 to (i) relaxing white-box access, and also (ii) to an adversary manipulating texts that affects both type I and type II errors.
For experiments on the convergence of the log-perplexity to the average string entropy, please see Appendix 5.
5.1 LLM attribution experiments
For our experiments in this section, we use two datasets that cover a variety of domains and LLM use-cases. We use news articles from the XSum dataset [44] to represent fake news detection, and Wikipedia paragraphs from SQuAD contexts [51] to represent LLM-generated academic essays. For details about datasets, please refer to Appendix B.2. To create LLM text, we first randomly sample 500 texts from each dataset. Then, for each sample, we prompt the first 30 tokens to each LLM and ask it to generate a 300-token text. The language models we use for generating the 300-token texts are LLaMA2 (7B) and LLaMA3 (8B) from the same family, Falcon (7B), and Qwen3 (32B) as a large model. We have white-box access to all models, which aligns with the Proposition 1’s requirement. In attribution experiments for each LLM, the evaluator LLM is the set $\mathcal{A}$, and the other three LLMs are the set $\mathcal{B}$. The null hypothesis is that the text is generated by one of the models in $\mathcal{B}$. Type I error occurs when the test incorrectly concludes that the text is generated by the evaluator model. Type II error occurs when the test fails to notice that the evaluator model generated the text.
Results in Table 1 demonstrate the high accuracy of our framework in the LLM attribution task. The LLM in the headline of each column stands for the evaluator LLM.
XSum (prefix = 300) SQuAD (prefix = 300)
Error Type LLaMA2 LLaMA3 Falcon Qwen3 LLaMA2 LLaMA3 Falcon Qwen3
Type I Error 0.02 0.00 0.02 0.00 0.02 0.00 0.00 0.02
Type II Error 0.00 0.14 0.00 0.00 0.00 0.10 0.04 0.00
Table 1: Type I and Type II error rates (rounded to 2 digits) for LLM attribution on XSum and SQuAD datasets.
5.2 Detection experiments
Setup & Datasets. To examine Proposition 2, we use human-written text to represent the null hypothesis (not LLM $A$). For human-written text, in addition to the Xsum and SQuAD datasets, we use prompted stories from the Reddit WritingPrompts dataset [13] to represent human creative writing. For details about datasets, please refer to Appendix B.2. We randomly sample 500 texts from each dataset, and prompt the first 30 tokens of each text to the LLM to create an LLM-generated version. The language models we examine are LLAMA 3 (8B), GPT-NEOX Erebus (20B), and QWEN (32B). We have white-box access to all models, which aligns with the Proposition 2’s requirement.
Test & Metrics. For this section, since the first 30 tokens are generated by a human for both null and alternative, we use the one-sided version of our test that removes the impact of the differences between cross-entropy and log perplexity caused by the initial 30 tokens. For a detailed explanation, please refer to Appendix B.3. Here, True Positive Rate (TPR) is the rate of correctly spotting that the evaluator model generated the text, and False Positive Rate (FPR) is the rate of incorrectly flagging human text as generated by the evaluator model. Note that FPR is equivalent to Type I error, and TPR is equivalent to 1-Type II error. Unlike Proposition 1, which is threshold-independent, Proposition 2 relies on a fixed threshold $t$ for rejecting the null hypothesis. Varying this threshold affects both the FPR and the TPR. Since different detectors use distinct methodologies and require their own thresholds, it is essential to account for the impact of these thresholds when evaluating the test performance. To address this, the literature typically adopts two approaches: (1) evaluating the TPR at a specified low FPR, e.g., [24, 11], or (2) using the Area Under the Receiver Operating Characteristic (AUROC) curve as a summary measure of performance, e.g., [66, 42]. Accordingly, we use two metrics. Our first metric is TPR, while FPR is held less than or equal to 1%. Note that the results of experiments in the literature, e.g., in [11], suggest that the FPR / type I error for current detectors can be dangerously high, which makes them unfit for regulatory applications. Our experiments using a TPR at a low FPR metric show the validity of our theoretical results and our method’s potential for regulatory application. Our second metric is AUROC, which plots TPR against FPR at all possible thresholds.
Benchmarks. We compare the performance of our method with various zero-shot methods. In choosing the benchmarks, we follow the literature, e.g., [42]. The log-probability-based benchmarks use the source model’s average token-wise log-probabilities. The (log-)rank-based benchmarks use the average observed (log-)rank of the tokens in the passage based on the model’s conditional distributions. The entropy-based benchmarks hypothesize that LLM texts are less-surprising for the model, leading to lower entropy predictive distributions. We also use Binocular [24] which has shown to substantially outperform all the zero shot detectors proposed in the literature in terms of TPR at low FPR.
Detection results. Table 2 shows TPR, while FPR is held less than or equal to 1%. The results show the higher performance of our method in fact-based datasets (e.g., SQuAD), where the human and LLM texts are less different, resulting in lower performance of the existing detection methods. Even when the FPR is held low, the human and LLM texts are more distinguishable in datasets with more creative content (e.g., WritingPrompts). Also, we observe that smaller language models (e.g., LLAMA 3 (8B) are easier to detect even with current detectors, and the maximum gap between ours and the second-best methods is in the SQuAD dataset with QWEN (32B) as the evaluator. Results for the AUROC are in Appendix B.1.
@@ 5.2 Detection experiments
5.3 Robustness experiments
Setup, Datasets & Metrics. In our theoretical results, we assumed white-box access to the generator models. In this section, we empirically evaluate the robustness of the statistical test in Proposition 2 under the black-box setting. As in section 5.2, we consider human text as representing the null hypothesis and text generated by a language model as representing the alternative hypothesis. We do not have access to the specific model that generated the text in the alternative set.
Instead, we leverage a standard assumption in the literature that language models are more similar to each other than to human text to design our test (e.g., see section 4 in 48). In the context of Proposition 2, this implies that the cross-entropy between human text and a language model is higher than the cross-entropy between two language models. We use the scoring model tiiuae/falcon-rw-1B to compute the log-perplexity and average entropy of text samples, and then apply the test as defined in Proposition 2. This evaluation setup is consistent with prior work on black-box detection (e.g., 42, 11). Following [11], we use TPR at 5% FPR as our evaluation metric.
We use the RAID dataset [11] that consists of 14,971 human-written texts from public pre-2022 datasets from 8 domains of abstracts, books, news, poetry, recipes, Reddit, reviews, and Wikipedia. For each human text, [11] creates a corresponding generation prompt using a template such as "Write a recipe for title". Then, they generate one output for each of the 11 models, 4 decoding strategies, and 11 adversarial attacks. The 11 models are: open-source chat models Llama-c, Mistral-c, and MPT-c; open-source non-chat models Mistral, MPT, and GPT2; closed-sourced chat models c-GPT, GPT4, and Cohere; and closed-sourced non-chat models Cohere and GPT3. They consider 4 decoding strategies: (i) greedy, which picks the highest probability token at each generation or sampling, which samples according to a generation probability $p_{n}^{B}(.)$, and (ii) repetition penalty or no repetition penalty. The repetition penalty evades detectors by reducing the likelihood of generating previously used tokens. This is achieved by down-weighting their probabilities by a factor $\theta\in\{1,1.2\}$, resulting in less repetitive output. After generating each text, they apply adversarial attacks that manipulate text on all samples, including human text and the $11\times 4$ LLM generations. These manipulations induce type I and type II errors in the statistical test. The nine query-free attacks are: Alternative Spelling (AS), Article Deletion (AD), Insert Paragraphs (IP), Number Swap (NS), Misspelling (MS), Synonym Swap (SS), Upper Lower Swap (ULS), Paraphrasing (PP), and Whitespace Addition (WSA). For details about decoding strategies and adversarial attacks, please refer to Appendix B.4. The described process yields a total of 134,739 human texts and 5,928,516 LLM texts. Given the cost of more than \$10,000 for re-running the experiment for the performance of benchmark detectors on approximately 6M texts, we use the results of [11]’s experiments on the other detectors and run the experiment for our own detector. For details about hardware and computation costs, refer to Appendix B.6.
@@ 5.3 Robustness experiments
Benchmarks. We compare our performance with multiple neural, zero-shot, and commercial detectors. The neural benchmarks are RoBERTa-B (GPT-2), RoBERTa-L (GPT-2), RoBERTa-B (ChatGPT), and RADAR. The zero-shot benchmarks are GLTR, Fast DetectGPT, Binoculars, and LLMDet. The commercial benchmarks are GPTZero, Originality, Winston, and ZeroGPT. For details about detectors, please refer to Appendix B.5. Recall that in this section, we do not have white-box access, and we use tiiuae/falcon-rw-1B as our scoring model.
Black-box detection results. Results in Table 3 demonstrate high performance of our detector in black-box detection with and without sampling and repetition penalty. For example, while other detectors have low TPR when detecting open-source non-chat models with sampling and repetition penalty, we achieve a TPR of 73.3%. Furthermore, we observe that, unlike all other zero-shot and neural-based methods, the performance of our framework never falls below the low threshold, with our lowest performance being 54.0%. The average performance of our detector in black-box detection is 82.9%, which (except for Originality, a commercial detector with TPR=85.2%) is higher than the average performance of other detectors, with the second highest performance being Binoculars with TPR=80.6 %.
@@
(llama-c, mistral-c, mpt-c) (mistral, mpt, gpt2) (c-gpt, gpt4, cohere) (cohere, gpt3)
Dec. Strategy greedy sampling greedy sampling greedy sampling greedy sampling
Rep. Penalty? $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
R-B GPT2 84.1 52.3 77.9 26.2 98.6 44.1 60.5 35.4 70.9 41.7 65.1 52.5
R-L GPT2 79.7 41.1 71.4 19.5 98.5 43.0 67.2 53.4 61.4 34.7 61.1 48.6
R-B CGPT 80.2 63.3 75.0 39.3 53.3 26.4 14.9 1.7 59.1 38.1 46.5 39.0
RADAR 88.8 77.4 85.6 66.4 91.8 63.8 48.3 31.8 81.6 75.3 72.2 67.7
GPTZero 98.8 93.7 98.4 82.5 74.7 34.6 9.4 4.8 92.3 88.5 60.6 53.4
Originality 98.6 86.3 97.7 72.5 99.9 64.1 89.0 51.2 96.8 89.0 91.7 85.4
Winston 97.2 90.1 96.6 78.3 68.2 49.0 29.5 11.3 96.1 93.7 73.2 68.1
ZeroGPT(*) 95.4 80.7 90.5 54.9 85.1 57.2 16.0 0.3 92.1 65.8 83.4 72.7
GLTR 89.8 67.5 83.9 38.3 99.6 56.9 44.5 0.5 80.7 54.3 75.6 63.7
F-DetectGPT 98.6 74.5 96.2 40.5 97.8 56.1 79.7 0.6 96.0 74.1 93.8 86.3
LLMDet 55.5 30.2 47.5 16.5 74.8 27.0 38.4 3.7 35.8 18.5 40.0 32.9
Binoculars 99.9 86.6 99.7 60.6 99.9 62.3 72.4 0.6 99.2 92.1 99.0 95.0
Ours 99.4 86.4 97.9 62.2 90.8 84.0 54.0 73.3 96.4 70.2 94.4 86.3
Table 3: TPR at FPR=5%. Cell colors: Red < 33, Yellow 33–66, Green $\geq$ 67.
Adversarial attacks results. Our detector maintains relatively high accuracy across multiple adversarial scenarios, with our lowest TPR at FPR=5% being 48.6%, which is higher than the lowest performance of all other zero-shot and neural-based detectors. Like other zero-shot detectors, our performance is reduced under Synonym Swap and Whitespace Addition, which is predictable because zero-shot methods often rely on fixed token distributions.
None AS AD IP NS MS SYN ULS WSA PP
RoB-B GPT2 59.1 55.6 37.1 56.9 55.9 43.8 71.5 18.8 45.2 68.9
RoB-L GPT2 56.7 52.4 33.2 55.1 51.7 39.5 79.4 19.3 40.1 72.9
RoB-B CGPT 44.8 43.3 38.0 5.2 44.3 42.1 39.6 31.7 0.1 49.2
RADAR 70.9 70.8 67.9 73.7 71.0 69.5 67.5 70.4 66.1 67.3
GPTZero 66.5 64.9 61.0 66.2 65.8 65.1 61.0 56.5 66.2 64.0
ZeroGPT 65.5 65.4 59.7 64.9 64.7 64.7 18.8 54.5 64.2 46.7
Originality 85.0 83.6 71.4 85.1 86.0 78.6 96.5 75.8 84.9 96.7
Winston 71.0 68.9 66.9 69.8 69.0 67.5 63.6 56.8 46.8 52.6
GLTR 62.6 61.2 52.1 61.4 59.9 59.8 31.2 48.1 45.8 47.2
F-DGPT 73.6 71.6 64.7 72.0 68.2 70.7 34.0 60.4 64.4 71.8
LLMDet 35.0 33.9 27.4 27.2 33.8 32.7 27.3 23.4 4.4 28.5
Binoculars 79.6 78.2 74.3 71.7 77.1 78.0 43.5 73.8 70.1 80.3
Ours 82.5 80.6 74.4 81.1 74.8 79.8 48.6 72.5 60.8 74.2
@@
Table 4: TPR @ FPR=5%. Abbreviations are: AS: Alternative Spelling, AD: Article Deletion, IP: Insert Paragraphs, NS: Number Swap, MS: Misspelling, SYN: Synonym Swap, ULS: Upper Lower Swap, WSA: Whitespace Addition, PP: Paraphrase. Cell colors: Red < 33, Yellow 33–66, Green $\geq$ 67.